Blog Image

Earthshine blog

"Earthshine blog"

A blog about a telescopic system at the Mauna Loa Observatory on Hawaii to determine terrestrial albedo by earthshine observations. Feasible thanks to sheer determination.

JD2456717

Observing log Posted on Feb 28, 2014 02:07PM

Tested a script while looking through the telescope. Script was set up to switch filters and take sequences of images.

Everything seems to be badly scrambled! Colours of filters do not match what is requested, and the sequences of images received are incorrect – for instance, when asking for 10 images at 2 seconds each I get about 3 images at 2 seconds and the rest are rushed through rapidly.

Shutter opens at strange times after a sequence.

It seems like the system is badly shaken up now. I think some of the above behaviour may explain some of the odd image-sequences we had at MLO, but the majority from MLO were good: now it all seems random.

My guess is that the LabView software has been jumbled somehow – it is easily done since the coding is not deterministic as text code is – it is all little coloured lines and boxes. It is easy to pick something up and drop it down in the wrong place. A test and a fix for this might be to restore an older version of the software – for instance something after Ingemar fixed the bad coding problems we found while at MLO and before the system was hit by lightning.

Another possibility is that hardware has been damaged in such a way as to scramble/ignore software instructions. Not sure how to test for this – we used a ‘sniffer cable’ to listen to the erroneous traffic to the mount, but I am not sure what to listen in on here?

Wonder if there is a way to ‘reset the hardware’? Perhaps the FW are ‘off by one’ – some latch is not clicking home?

—————

Inside the ‘breakout box’ there are lights to watch as a script runs. There is a light on a ‘hager’ relay and it evidently lights up when the two Thorlabs filtwer-wheel controllers are being adressed. They have displays that hsow numbers and I am guessing they show the position of the FW.

While running a sequence through the 5 filters I saw the numbers change. I think I caught a rythm corresponding to several passes through the 5 filters, but I swear there was an irregularity [hard to prove afterwards as there is lots of uncomfortable crouching and waiting to do]. Also strange is that the FW displays show two sets of numbers first, these then change to another set, then the hager light goes out and you can hear the shutter start working [its progress is ueven, however, despite being asked to take 10 identical images]. Here is a sequence:

FW2:5->5,6->1,2->3,4->2,3->4,
FW1:2->1,2->1,2->1,2->1,2->1,

where “A->B” means that the FW display first showed an A which then changed to a B, after 5 secodsn or so. Notice how FW1 alwasy goes “2->1”. FW2 seems to be cycling all numbers. Notice how the start position in one par is alwasy 1 highe rthan the end position in the previous set. There is a cryptic sentence deep insdie Dave’s LabView code about how Thorlabs uses one set of indices and his code another – this coul dbe it, I guess.

I think FW1 is the ND filter and is always set the same way – in the above test I was asking for the ‘AIR’ position and I think this is what we see. I will now run a test where I ask for one of the ND filters instead and we will see if the numbers change!

—————

OK, I did that, and it seems I am right. FW1 is the ND filter. I set it to ND2.0 and the sequence 2->1 above changed to 5->6.

Should run a sequence now of just changing the ND filters.

——————

Ok, did that too. This time I used a scriptthat always set the ‘B’ filter, but cycled through the available ND grades – that is AIR, ND0.9, ND1.0, ND1.3, ND2.0. The typical sequence of readouts from the Thorlabs controllers was:

FW2:3->2,3->2,3->2,3->2,3->2, …
FW1:6->1,2->1,2->3,4->4,5->5, …

we see how the colour filterwheel (FW2) always sets the same filter and always seems to start at the one numbered one higher. Visual inspection during the cycling revealed that the colour of the light switched between the colour of the lamp and Green – i.e. the FW2 settings actually acquired were ‘Air’ and ‘V’ – not ‘B’ as I asked for.

FW1 meanwhile seems to have cycled irregularly – 1,1,3,4,5, … (2 is missing). It was not really easy to see whether various ND grades were inserted or not, but clearly AIR was one of them since I could see colours of the light.

Note that in a sequence “A->B” a sound – such as of a filter-wheel changing position – is heard before A, and before B.

Not sure what to make of this. It could be a signals problem where feedback from the actual position of the wheel is missed or missunderstood by the software so that new or wrong commands are sent.

Is the software that directs the FW giving absolute commands or relative ones, as in
“Go to the blue filter”
vs
“go to the filter after the one you are at, which I assume is the IRCUT filter, so that you will arrive at the B filter since it comes after IRCUT”?

Why does the wheel change position twice during a set the FW-wheel command? Makes sense only if it always went to some home position, and then proceeded from there – but clearly it is not starting from the same position each time, given the readout above. Again, could it be that the intent is to send it to home each time but this goes wrong and it doesn’t actually arrive at home, but assumes it did for the next relative command?

NEXT time, we need ABSOLUTE ENCODERS. I want a little man inside the telescope with a cell phone calling us with the ACTUAL position of each FW!



JD2456715

Observing log Posted on Feb 26, 2014 02:02PM

Manually testing the shutters and FWs.

We disconnected the CCD camera so we could look through the telescope along the optical axis, at a lamp. We ran the camrea so that the internal shutter opened and closed regularly, and then manually changed the color and ND filters.

Strange behaviour was observed: Setting the color of the filter was not a simple affair – the colour would be a bit random, and the ND filter would reset – apparently by itself – after a few seconds.

We will try to run scripts next, to see if the command of the devices from the scripts are more consequential and whether the ND resets in the middle of a sequence. Our experience from masses of observations from MLO (before the lightning struck) is that stacks of images are all OK as long as the first image is all right.



Shutter/FW failure – 2456693.2

Observing log Posted on Feb 04, 2014 04:31PM

Testing the system on a lamp-illuminated flat field. With a script, sequences of images are taken with commands to change filters. Kinetic sequences.

There is obviously some problem with either the shutter or the setting of the FW, because while some images are well exposed at more or less recognizable exposure times (we use a different lamp from the hohlraum we had on MLO) some are simply almost all black at those same exposure times.

To test if it is the shutter that is failing to open, or the FW that does not set, we must eliminate one thing at the time – first let us run a script with exposure times of varying lengths, and simply listen for how long the shutter is open.

If it seems the shutter is always open as long as we ask it to be, we move on to the FW – perhaps we can open the telescope box and simply look at whether it turns (pant a white spot on the side!).



More progress … and yet …

Observing log Posted on Jan 23, 2014 11:24AM

As stated before, scripts can now be run and images gathered. We have solved a minor problem with disc-access that suddenly popped up. Still not sure what this is due to, but can now save images from automatic scripts. This clears the way for a row of lab tests that will start now.



Scripts running

Observing log Posted on Jan 21, 2014 12:40PM

It was possible to start a simple observing script today: Mount was not attached so nothing moved except shutters and FWs (and the SKE is still not working) – files were obtained but it turned out that something with the setup of permissions on the PXI has changed so images could not be written to the ‘NAS’ drive (a network drive in the instrument rack). Am seeking help from DMI IT department for this.

If that starts working we can run tests on bias fields at least, and later on flat fields (with a screen of some sort in the distance) and then focus tests and all this.



CCD is back

Observing log Posted on Jan 17, 2014 04:25PM

We managed to get actual images from the CCD camera today. Hans will tell the story of what was wrong.

So – testing all systems together, next!



… and Dutch Radio!

Observing log Posted on Jan 15, 2014 03:36PM

‘de Kennis van nu’. The program will be available online, from the web-link http://www.wetenschap24.nl/programmas/de-kennis-van-nu.html



DMI news item

Observing log Posted on Jan 14, 2014 04:34PM

Our paper also made it to an online DMI news item.
http://bit.ly/1eBN5K2



Status JD2456645

Observing log Posted on Dec 18, 2013 10:30AM

Sloooowly the system is coming back to us: We now have LabView Engineering Mode control of Front shutter, Iris, Mount, Filter Wheels, and now also Focus stage.

We still need to get the SKE stages working.
We need to plug in the camera and see if it can be brought under LV control.

With these things running we are closer to getting some technical data from the system, such as bias frames, all-colour flat fields (for the first time), and we could study the shape of the PSF in different filters.

Important next steps are:

a) get the CCD working and under LV control
b) Try running simple scripts again without telescope mounted

Future, ambitious, steps are:

c) assemble the actual telescope in a testing facility and run pointing tests – perhaps test long sequences of observing scripts indoors and actually debug the system so it would work under field conditions.
d) See to the never-attempted integration with a weather station



More small steps

Observing log Posted on Dec 05, 2013 02:17PM

Got the FWs moving today, using very basic LabView modules. Engineering Mode does not seem to want to talk to its ‘subVIs’ any more.

For reference:

in the VI ‘set DO line’ you can select port and line combinations. So far we know that:

port0/line22 operates the realys in the Breakout Box labelled RE2 and RE4.
port0/line23 – RE3 och RE5
port0/line20 controls the Thor FWs. The two relays adjacent to the Thor controllers come alive. The FW can be selected via the pushbuttons on the boxes.

If we can get the camera alive next we could at the very least do lab work on the CCD properties – make color flat fields, study the PSF in various filters, and so on.



Small steps …

Observing log Posted on Dec 04, 2013 04:21PM

Today Hans was able to operate the front shutter and the front iris using LabView. The licensing problem seems to have been cleared up (it seems to have been a typo in a form somewhere – perhaps it happened at MLO, or later at DMI). The CCD camera and the Magma no longer are listed with ‘yellow triangles’ in the system – we have yet to plug it in and see it operate.

Next steps are to use same methods to get all devices running – Engineeering mode seems to be dead somehow, but all the ‘sub VIs’ are there and so far we have gotten the front shutter and iris VIs to work – can slew telescope from sub VI also.



LabView code

Observing log Posted on Aug 02, 2013 12:29PM

Ingemar was here and looked at the LV code. We semeed to learn the following:

1) The NI Report Generation Toolkit license has run out and needs replacement.

2) The newest set of LV codes are on the NAS under ‘Earthshine Project/LabView Sou…/’

3) The Engineering mode refuses to run because of some error messages related to the AXIS – these are for the FWs, the SKE and the Dome. No matter what you want to do with the Engineering Mode it will always start by initializing all its sensors, and if one of them fails, nothing works! Since we have no dome there will indeed be a problem here!
Ingemar was able to turn off some of this (NB: set it back to as-was, later!) and indeed the code progressed to a further point (when it then ran into other problems). Front shutter was tested while the error-checking was turned off in Eng. Mod. but did not work.

Of the above 1) may not be important as long as scripst are not being read. But 3) seems pretty serious.

Questions:

Are the new licenses for the MAX correctly installed?

Can the code that refers to the dome be turned off?

Will things work again once the dome is turned off and the other AXIS cables are plugged in?

Why did front shutter also not work – it has no motor?

Added later: Some of the above is now clearer. See here. A number of the errors were related to wrong LV license number, and the new IP numbers for DMI must be set insid ethe LV code. For reasons we do not understand the COM-port numbers had changed.



Data inventory – non-lunar images

Observing log Posted on Jun 06, 2013 10:23AM

Here is a list of non-lunar images taken at MLO. There are stars, clusters, galaxies, and planets and asteroids. Many are suitable for studies of extinction.



Best of the data — averaged frames with 100 good slices

Observing log Posted on Dec 14, 2012 05:58AM

I wrote a program called moonraker in fortran, which processes all our data, seeking out the best for further reduction. Peter has written something similar in IDL.

It goes through all the images and throws out ones with

1) saturated BS (> 50,000 counts in the peak of the BS)

2) underexposed BS (<10,000 counts)

3) smeared images (heavy bleeding of photons in the Y-direction on the chip, reaching the bottom 20 pixel wide strips of the frame). A mean count level of more than 50 counts in this strip triggers a flag and the frame is rejected.

4) rough check if the moon is too close to the edges of the frame. This is not yet optimal, but throws out the worst cases.

5) checks that the temperature of the CCD is OK (an entry in the fits file header tells us if the temperature of the CCD has stabilised or not)

If all 100 slices in multiple exposure frames are OK, according to the above, they are bias subtracted (using the biases on either side of the exposure, averaged appropriately for each slice, and scaled to the superbias created by Henriette) and averaged, and written to disk.

6) The magnitude of the total light in the frame (i.e. apparent magnitude) is then compared to the expected V magnitude (using the JPL model, with corrections for the east and western sides of the moon, depending on which side is illuminated by BS), taking into account the airmass of the observation and the extinction for the particular filter. This isolates for removal exposures which have been taken with an incorrectly reported filter. Finally, after extinction correction, only those frames which fall within ~ 0.2 magnitudes of the correct magnitude using the JPL model are retained.

The list of the final set of good exposures using this method is attached (all100.doc — actually an ascii text file):

There are 534 good frames, spread from nights JD2455938 to JD2456104.

A summary is as follows:

Night, number of good frames in V, B, VE1, VE2, IRCUT, comments

JD2455938 2 2 0 5 0 fullish moon, some foggy frames
JD2455940 3 1 0 2 0 fullish moon
JD2455943 3 3 0 3 1 fullish moon
JD2455944 3 1 2 1 3
JD2455945 1 1 1 2 0
JD2456000 3 0 0 1 1
JD2456002 7 0 0 0 0 V band only
JD2456003 9 0 0 0 0 V band only
JD2456004 1 0 0 1 0
JD2456005 6 0 0 0 0 V band only
JD2456006 4 0 0 0 0 V band only
JD2456014 3 0 0 0 0 V band only
JD2456015 7 10 4 6 4 1/3 moon, lots of frames
JD2456016 11 10 9 11 11 “
JD2456017 7 11 0 11 2 “
JD2456028 5 5 0 9 0
JD2456029 0 0 0 6 2 2/3 moon, lots of scattered light
JD2456030 1 0 0 3 2
JD2456032 1 0 0 4 1
JD2456033 4 1 4 2 4
JD2456035 0 0 0 1 0
JD2456045 5 5 4 4 4
JD2456046 4 8 10 11 10
JD2456047 6 5 8 6 7 stars visible in some frames, useful for PSF
JD2456061 2 4 3 6 4
JD2456062 0 0 0 2 0
JD2456063 0 0 0 5 0
JD2456064 0 0 0 3 0
JD2456073 4 7 7 10 8 some eclipsing of moon – dome issues?
JD2456074 4 2 3 7 5 scattered light haze or fog?
JD2456075 8 7 8 7 8
JD2456076 9 1 0 2 0
JD2456089 3 2 1 1 0
JD2456091 2 1 5 0 4
JD2456092 0 0 3 0 5
JD2456093 0 0 1 0 5
JD2456104 5 7 5 1 5 some foggy frames, moon a bit close to bottom corner

I’ve inspected all of these by eye and they look mostly pretty good! There are still a few bad frames here because the scattering of the halo is not selected for yet — but the number of frames with low alpha (i.e. haze or fog) is just a handful.

Some of these nights will be pretty useful for looking at colour changes in the ES over time. I am doing that next.



JD2456109 – weather log

Observing log Posted on Jul 01, 2012 12:06PM

Clear. Moon up.
7-9 m/s from 160 deg.
14% RH rising
7 deg C.



JD2456105

Observing log Posted on Jun 27, 2012 01:15PM

I do not understand this image:

The image is not overexposed (3000 counts only) so the streaks are not ‘blooming’, which occurs when the pixels are saturated (starts near 60000 counts). A streak downwards would be due to readout while shutter was open (down corresponds to the readout direction). But UP is very strange – it could happen if the telescope was moving while the Moon was being imaged, but as far as I know this is not a situation we have seen before. It could also happen if the readout direction changed – but I do not think that is possible – and in this case readout would have had to occur in both directions then.



JD2456102

Observing log Posted on Jun 24, 2012 01:52PM

Very clear. High wind.

Shutter failed pretty miserably – 1 image stack of 16 was useful.

Later restarted telescope and was able to calibrate the mount – this means that the shutter started working again.

Google Sky helps find the calibration position.



JD2456095 – weather log

Observing log Posted on Jun 16, 2012 03:59PM

Hazy – VYSOS line is yellow/white. Moon rises in clouds.

8 m/s from 180 deg
RH is 15%



JD2456093 – weather log

Observing log Posted on Jun 14, 2012 04:11PM

clear, 40F, 15-20% RH rising, 6-8mph or 2-4 m/s from 220 degr.

Many ‘hung shutter’ frames but the ones that work are perfect.



JD2456092 – weather log

Observing log Posted on Jun 13, 2012 03:25PM

clear, 35 F, 30% RH, 2 mph winds.



JD2456090 – weather log

Observing log Posted on Jun 12, 2012 02:18PM

hazy
40 F
RH is 30% and rising
4-6 mph winds.



JD2456090 – weather log

Observing log Posted on Jun 11, 2012 03:14PM

9 C/50 F, 8mph/3-4 m/s, 15% RH, clear but not v. so.



Data status May 2012

Observing log Posted on May 30, 2012 02:05PM

As of end of May 2012 we have about 25000 images labeled “MOON” – many of these contain 100 images each.

From these, after coadding all image-stacks, and by selecting only good data, we get 6871 single images (these are now either single observations or aligned 100-image bundles). These cover 55 separate nights from July 2011 to now.

Most of these 55 nights are observed in all filters.

None are science-grade SKE or ND images.

With an expected 200 good nights and a duty cycle of 20-25% we are not doing too badly!



JD2456076

Observing log Posted on May 29, 2012 08:41AM


Shutter all over the place today – dragging AND overexposure??

9-11 m/s from 120 deg. RH near 20%, 6 deg C. clarity 9/10.



JD2456075 – weather log

Observing log Posted on May 28, 2012 08:11AM

Hazy or foggy – quite large halo on Moon in TV cam. Difference image showing structures.

15 m/s from 160 deg, steady. 6 degrees C. RH is 30%.



JD2456074 – weather log

Observing log Posted on May 27, 2012 07:27AM

12-18 m/s from 140 deg. humid – RH is 70% and rising. T is 5 deg C. It is almost clear – but minutes ago small puffy clouds blew over …

Ups — more clouds. Closing down for a while now.

Observing again – must say there is some fog banks that come and go. Not one of the best nights.



JD2456073

Observing log Posted on May 26, 2012 09:45AM

Clear. 12 m/s from 150 deg. 48 deg F. 10% RH.

At 2456073.8ish the cables are in the frame.

Good data – butthe usual filter dragging and blanks. Especially V which is unusual.



JDS2456071 – weather log

Observing log Posted on May 25, 2012 08:09AM

18 m/s from 160 deg. RH 20%. 5 deg C. Clear.



JD2456070 – weather log

Observing log Posted on May 23, 2012 10:57AM

Clear. 15 m/s from 140 deg. RH is 20%. 4 deg C.



JD2456069 – weather log

Observing log Posted on May 22, 2012 08:23AM

Windy at 17 m/s from 160 deg. Clear. RH 50%. T near 4 C.

Focus problems – again!



JD2456063/4

Observing log Posted on May 16, 2012 09:41AM

Trying to determine if CCD is linear by observing pairs of stars at various exposure times. Finding that focus is bad !?Is it even possible to get ‘doughnut’ images for out-of-focus stars in a refractor?RH is 50%, T is about 44 F, wind is 6-8 mph- It seems clear.Later, osberving Moon: ..54.06.. T=37ish F, RH is down to 42%, wind 2-3 mph. seemed clear when the Moon rose.



JD2456062/3 – weather log

Observing log Posted on May 15, 2012 10:31AM

at ..62.8 clear, 3 deg C, 2 m/s from 250 deg. RH 40-55%.at ..63.11 T lower, still clear, low wind, RH 60%This is a really strange night – the relative exposure times are very unusual.The factors are:B 20V 22VE1 20VE2 18IRCUT 20Especially B is very strange indeed.A lot of shutter hang going on, though.



JD2456061/2 – weather log

Observing log Posted on May 14, 2012 12:49PM

3 m/s from 220 deg, 4 deg C, RH 35% and rising.
Zenith clear but Moon rose behind clouds.

When observing resumed there was a slighthaze – VYSOS cloud line sloooowly moving upwards as dawn approached – but at least no obscuring clouds, as was the case at Moonrise!



JD2456060 – weather log

Observing log Posted on May 13, 2012 10:34AM

Clear
T about 4 C
RH down to 15% from a high hours earlier.
3 m/s winds



JD2456046 – weather log

Observing log Posted on Apr 29, 2012 08:04AM

12.8 degrees C in dome at start – 45 F outside on MLO mast
9 m/s increasing at start. At xxx.86 the wind is 16 ms or 20 mph. Steady from South.
RH 15%
clear

Exposure factors for the filters are very good this night:

B 34.0
V 11.8
VE1 2.6
VE2 10.0
iIRCUT 2.6

These are factors, note xposure times. The exposure time is some number times the above. A typical IRCUT exposure time may be 0.008 s. The code we have – script_builder_MOON.pro – scales a table like the above one appropriately for the lunar phase.

Later: The above factors seem inadequate on another night with the Moon rising – B is overexposed while VE1, VE2 and IRCUT are at 12000. Hmmm.

A bit later still: Hmmm – the above problem was due to hanging shutter – but a very consistently hanging shutter! So the table is OK. B is perhaps a bit high but now V and VE2 are just right.



JD2456045 – weather log

Observing log Posted on Apr 28, 2012 08:24AM

12.9 C in dome at start (6045.75..)
RH 25%
0 wind
clear
Moon setting

Near ..45.86.. one of the cables from the MLO tower gets in the image.



JD2456034 log entry

Observing log Posted on Apr 16, 2012 03:56PM

Moon rising in a cloud.
50% RH.
Temp=35 F.
11 mph winds.



JD2456006

Observing log Posted on Mar 19, 2012 05:12PM


Extraordinary night – followed the Moon from moonrise until dawn. Clearest night we have had yet. Halo almost non-existent on DS.

This shows that we may be able to get very good data very close to the Sun – but in that case extreme refraction would have to be handled. Will look into whether this can be achieved from the JPL ephemeris. Exposure time also needs to compensate for the strong extinction through 5-10 airmasses. Not sure how to do that without stopping and starting scripts.



Data inventory March 2012

Observing log Posted on Mar 19, 2012 01:43PM

As of March 18 2012 we have 78.000 images of the Moon, of which 51.510 are relevant for data-reduction with the current techniques at hand.

Distribution across lunar phases is seen here:



JD2455996

Observing log Posted on Mar 09, 2012 10:05AM

Hazy – ring around the Moon.



Wild Slew is back!

Observing log Posted on Mar 07, 2012 08:21AM

On JD2455993 I tried to do lamp flats since the night was cloudy and humid. Started the LAMPTEST script, the system started slewing towards the hohlraum lamp – and kept going towards the floor. KS stopped it before it hit anything.

Why on earth are we back to the ‘old ways’? This is the sort of behavior we had before the ‘dont send commands during slewing’ problem was solved by Ingemar.

Because the system has been well behaved for a long time I did not have the cable sniffer running so we have no record of what went on.

Any ideas of automatic operations while the system still does this must be ruled out!

Will now reboot everything, calibrate (on another night), and see if the problem reappears.

We can only hope that someone had been moving the telescope about by hand, as that would explain what happened.

T=7,4 C in dome, RH=70%.



JD2455991

Observing log Posted on Mar 04, 2012 09:29AM

Moon in clouds. RH 85%. No observing.



JD2455990

Observing log Posted on Mar 03, 2012 05:04PM

A slight haze at MLO.



JD2455988

Observing log Posted on Mar 02, 2012 08:15AM

Trying to observe Moon. Many thin clouds, but it may clear up later and then we will not waste time on setup. T=7.4 C. RH near 40%, varying.

Shutter works irregularly at first, then regularly – as if it has some dirt or oil in it that needs to be rubbed around until it is warm? Now and then there is ‘dragging’ but, as always, mainly in one filter – when a new filter is chosen the dragging may go away, or reappear in another filter.

Dome needed a fresh ‘find home’ – wonder why? I left it in a working condition? Dome now follows telescope.

Moon in frame but not well centred. Why? It is just a few days since I calibrated the mount on a target in the same part of the sky? ‘Slow creep’ of alignment seems to be a feature of this system.

More clouds now. Shutting down. Need a cloud sensor up and running!



JD2455987

Observing log Posted on Mar 01, 2012 08:12AM

Ring around the Moon! No observations! Doing lampflats instead. T=7.3 C in dome.



JD2455983

Observing log Posted on Feb 26, 2012 10:42AM

Trying to observe M41. Temperature in dome 6.4 C at start. Very light haze.

Shutter giving problems: So far, all shots are bias frames – except one that is 0’s in one third … as if the camera died?

At one point, when this happened, some LabView panels having to do with Andor configuration and Shutdown moved to the front on the logmein display: I think this means they had been activated – I had done nothing to cause this – so perhaps it means that the system did it having sensed something was wrong?

Abandoned M41 – then tried the hohlraum source. No images at all recorded now.

Rebooted everything.



JD2455982

Observing log Posted on Feb 25, 2012 01:14PM

Mount re-calibrated after Ben’s fix of the Kill Switch ring. Can now reach -30 degrees declination.



JD2455949

Observing log Posted on Jan 22, 2012 05:21PM

Ran scripts to observe Saturn and Mars. Strange problems with script not writing to disc. Rebooted and all was OK.

T 4C outside, 6.7 inside. 10 m/s, 40% RH. Clear.



JD2455944

Observing log Posted on Jan 18, 2012 11:38AM

Observing asteroid EROS since it is close to Earth. Trying for a movie of the Moon rising behind the flank of Mauna Loa.

45% RH (rising), some indication of clouds, T near 3 C outside, winds 4 m/s 220 deg.



JD2455940

Observing log Posted on Jan 13, 2012 05:37PM

T fell to -3 degrees C outside. Inside the dome it was as low as 2.7 degrees C. About 30% of images were all right. RH very low. Wind near 8m/s.

Hit kill switch again at Alt 75 degrees near meridian. Mount calibration may be lost now – but seems to be less than a few pixels.



JD2455938

Observing log Posted on Jan 12, 2012 11:38AM

Moon observing with script Moonb.csv, which – wait for it – AUTOMATICALLY adjusts the dome position! Whohoo! AND the exposure time!

T near 4 C in dome so many dragged frames. RH v. low. Winds near 5 mph. Halo around Moon seems large.

Good to listen to the shutters etc via a Skype call from the telescope. Shutter is LOUD, when it works.

Mount is re-calibrated following yesterdays chrash into the kill-switch.

Image still jumps, but not as much as before – it seems to be at only the 5-10 pixel level now.



2455933

Observing log Posted on Jan 07, 2012 10:31AM

Testing focus and testing the script MOONb.csv that should automatically set exposure time (software sets it automatically as it builds the script).

T 4 deg C, wind 5 m/s, RH 20%. Almost Full Moon. Slewing went well. Moon nicely centred.



Data inventory as of end of 2011

Observing log Posted on Dec 31, 2011 12:14PM

Here is a histogram showing the distribution of our observations across Sun-Earth-Moon angle (i.e. lunar phase on 0 to 180 degree interval). The huge spike near angle 50 is from the tau Tauri observations.



JD2455924

Observing log Posted on Dec 29, 2011 09:23AM

At sundown T=8.8 wind dropping, 10-14 m/s, RH about 20%. 2 hours later outside T was 5 degeres C while dome T was still 7.

Shutter failing often – and ‘open while reading out’ occurring. Again we see sets of images failing. Of course not related to the physical filter itself – but then why e.g. all VE1 images failing while adjacent images OK?

No clouds. halo starting to reach DS edge towards sky.

Noted that an antenna gets in the way of the Moon during some Moonsets.



JD2455923

Observing log Posted on Dec 28, 2011 05:47AM

Catching the setting crescent Moon. T=8.4 C in the dome. Got some nice Moon images – almost no scattered light!

Then observed Uranus, but images were either saturated or misfired. T=4.7 C now.

Now M41. Getting very few frames. T=4 C and the wind is strong at gusting to 16/17 m/s.



When fog falls yea verily upon the mountain

Observing log Posted on Dec 21, 2011 03:26PM

A skycam image showing where there is fog on the mountain. The navigation beacons are hazed out (in the NNW) so there’s fog at ground level. Close up please!



JD2455905

Observing log Posted on Dec 12, 2011 11:49AM

Dec 10 2011 Lunar Eclipse.Observed with Chris. Fairly clear skies. Temperature was sub 30 degrees F so plenty of shutter stick. Of 1500 images about 2-300 were useable.Made animated GIF of the eclipse.Image position wuite stable, but not perfect – clearly the fix by Ingemar which now enables interpolation helped a lot, but table may still be to coarse – otherwise slight ‘jitter’ in position may be due to ssome mount issue.



JD2455886

Observing log Posted on Nov 20, 2011 01:53PM

Clear, but 80% RH and T=5.6 C in dome.

Halo enormous – no sign of DS despite phase. Exposures of the BS only in the 1000’s – must update exposure time – heavy clouds??

Shutter clearly open during readout – dragging downwards.

Also sticking so that saturation in stripe downwards happens – i.e. shutter started to close, then stuck open for a long time. T now at 5.2 C.

T now at 4.9 C.

Experimenting with SETIMAGEREF and MOVEMOONTOREF because it seems the mount calibratio I did earlier tonight was not quite right. Noticing that when Moon is half out of the image frame downwards I get a reflecting coming in from the upper side of the frame – like was saw during the tau Tauri occultations in July – then it was left to right. So, something in the optical train is causing reflections onto the CCD when the source is not well centered.

More and more clouds now. T now 4.8 C.



JD2455883 darks

Observing log Posted on Nov 18, 2011 04:02PM

I have obtained 100 dark frames with random exposure times in the interval 10-200 sec. This will allow me to determine with certainty if there is any substantial dark current.

The telescope performed as it should and all images were saved without trouble.



Wet

Observing log Posted on Nov 18, 2011 11:06AM

Extremely wet and cloudy these days! Wanted to observe Moon.



JD2455877

Observing log Posted on Nov 12, 2011 10:19AM

Testing the Diffuser. it sits in the ND filter wheel – that is, in the collimated beam. Being situated there it removes any spatial information originating in the flat screen itself, the front objective and the first lens of the collimating system. Flat fields taken this way have a central brightening that looks spherical or parabolic in shape.

Perhaps, if that darkening could be understood and subtracted by some fitting procedure, we can use the diffuser to get good high spatial frequency information about the FF? It seems to remove the huge gradients that otherwise characterize the FFs we have been getting from sky, dome and lamp in the past.

The difference between darkest corner and centre is about 10%. What is the origin of the circular feature, and what does its position off-center tell us?

We fit a second-order surface to the image above and subtract it. Add the mean back in, and normalize. The result is here:

While extremely similar in appearance to the other ‘superflats’ we find that the central red patch is about 1% above the rest.



Refraction on JD2455729

Observing log Posted on Nov 01, 2011 10:51AM

On the July 26 2011 occasion of observing the star tau Tauri being occulted by the Moon we followed the Moon from just after Moonrise until Sun rose some hours later. During that time we tracked the star at sidereal rate so the Moon drifted through the frame.

Determining the radius and image-plane coordinates of the lunar disc we plot these:
We note that the radius increases slightly as the Moon rises – probably an effect of differential refraction decreasing as airmass approached 1. At most the radius estimate increased by about 1.5 pixels during the sequence. The standard deviation about the line is 1 pixel, telling us what the precision is of the method used to estimate lunar disc properties (it works by fitting a circle to points on the edge of the lunar disc).

We also see that the Moon did not follow a straight line across the image plane. The deviation is up to 5 or 6 pixels – about half an arc minute. This may be a refraction effect or a tracking effect or indeed an orbital effect – the mount we have does not track the Moon in declination.

In summary we see that the effect of differential refraction across the field is not more than half an arc minute at large air masses, consistent with physical estimates.



JD2455865

Observing log Posted on Oct 31, 2011 06:12AM

Observing Moon. Temperature in the 7-5 degrees C range (falling). Shutter sticking now and then. Weather very similar to previous days – clouds and rain into the late afternoon and then just at sunset the skies clear.

Dome ‘nudge’ working well and Dome Match also, although some work is needed on the table of pointings.

Halo around the Moon.



JD2455864

Observing log Posted on Oct 30, 2011 06:43AM

Getting the setting Moon in clear skies.

Plenty of ‘filter stick’ and it comes in sets – when we change filters the shutter ‘unsticks’ and we are good for one set of images, then we change filter and there is the chance that we get good data or bad data.

We also see a periodic shifting of the Moon inside the frame. Either the JPL table of Moon positions is slightly bad, or the interpolation is done.

Question: Is the table of Moon positions interpolated or do we go for ‘nearest neighbour’?


See the telescope and the Moon!



HD2455859

Observing log Posted on Oct 24, 2011 05:04PM

Strange ‘dragging’ on at least the VE1 images – and yes, readout drag is up-down, not right-left.

Many frame dropouts. Temp. is between 4 and 5 degrees C.

B and V perform well at this very almost new moon, but VE1 and VE2 seem to need boosting relative to the bluer bands.

Wait, the VE2 is fine – it is the VE1 that is too low, as if it, in particular, wants to stick.

Noting that the external view of the VYSOS-20 position is the best for detecting the coming of Dawn.

VERY little scattered light from the BS tonight – even at very high airmass.



JD2455858

Observing log Posted on Oct 23, 2011 05:39PM

4-5 degrees C. Less thin cloud than previous nights. Shutter misses quite a few images. Towards dawn shutter working better. Sky noticeably brighter by 5:25 on all-sky camera, but not in CCD images.



JD2455857

Observing log Posted on Oct 22, 2011 05:01PM

Getting multi-band images of crescent Moon. Bands of thin clouds rolling past – vissible mainly as increased halo around Moon and in the eq. diff. images. T between 6 and 7 degrees C. Not very many failed frames.

Dawn started 5:45. by 5:55 sky-level was 445 – normally 397.

Time Level
6:00 505
6:02 530
6:04 930 – bias pattern starts to be seen
6:11 Earthshine side weaker than patterns in bias. Stopped.



JD2455856

Observing log Posted on Oct 22, 2011 10:37AM

Sky flats in the morning. Stopped when trailed stars were seen. VE1 flats ‘blotchy’ and strange.



JD2455856

Observing log Posted on Oct 21, 2011 04:11PM

Lots of thin cirrus! 9 degrees C. Taking lunar images. Will follow the Moon up in case it clears before dawn.

Can confirm that GOTOMOON script command works – goes right to the Moon. [Added later:] but, weirdly, seems to drift slightly. Since the table is OK it must be the calibration that goes slightly off – polar axis alignment issue?

DOMEAZ was supposed to put the dome where the telescope is pointing – and it worked for a while, and then started sending the dome off to strange places. Perhaps some homing is needed on the dome before using the command?

Testing the combination

GOTOMOON
MOVEMOONTOREF

to see if it keeps the Moon centered. At least the MOVEMOONTOREF does not send the telescope anywhere crazy.


Nice screendump of what the Front Camera Saw earlier today. The Pleiades!



JD2455854

Observing log Posted on Oct 20, 2011 06:24PM

9 Degrees C. Some thin cloud from the start – then more and more.

Lots of Moon observations in ‘dither’ mode, through all filters.

Day started with CCD flatlining. Reboot. Then all OK, Still have to set sub-degree offsets on pointing to put Moon in middle. Why? Will test the MOVETOREF command.

FW apparently OK – no noticeable ‘sticks’.



JD2455853

Observing log Posted on Oct 19, 2011 02:00PM

Various things failed and none of the planned observations could be obtained:

Telescope did maverick meridian flip that we cut short as it neared the floor.

The camera flatlined for none of the known reasons.

Then some part of the FW or rotary stage system got stuck. There is an issue with the large rotary stage acknowledged by several parties. I additionally think the FW can get stuck at times – witness the failed VE1 observations on JD 2455849.

PXI is at the moment unreachable over the internet following an iBoot power cycle that failed to re-boot the machine.



JD2455849

Observing log Posted on Oct 14, 2011 06:37PM

More ‘dithered’ observations of the Moon. VE1 filter images kept being ‘bias frames’. How does the shutter know that the VE1 filter is in? Or is something else going on?

Pointing with absolute coordinates is awkward – from night to night the mount calibration seems to go slightly off so that new offsets have to be tested in a time-consuming manner. If the polar axis alignment is not good, would it cause this problem?

CCD seemed to show new startling ‘saturation’ modes.

here



JD2455838

Observing log Posted on Oct 04, 2011 07:51PM

Apparently managed to ‘unstick’ the previously stuck filter-wheel stage that caused blocking of the FOV by the edge of the aperture of the stage that holds the SKEs. This was done from the Engineering Mode by ‘finding the homes’. The Large Rotary Stage did not reset in such a way that the green light indicating success ever lit up. The Little Rotary Stage rotated, cleared the blockage, and its light lit up.



What the Skycam saw

Observing log Posted on Oct 03, 2011 08:29AM

A selection of Skycam shots to show what you can see. From top left : a hazy night, a very hazy night (ice crystals?), a very clear night, water on the sky cam and the laser star, the skycam on a wet afternoon and the crescent moon on a bright sky just before sunset.



JD2455836

Observing log Posted on Oct 02, 2011 08:56AM

Skyflats, Moon and Altair.

Skyflats started too late and are not well exposed – I started at 6:13, Sunset was close to 6:20 so should perhaps start closer to 6:00 next time.

Moon was run through various settings of SKEs and the KEDFs as well as the NDs – but images appear strange – as if the devices did not set right or the input data was wrong.

Altair was imaged through variable cloudiness allowing us to see both very clear PSFs and the misty PSFs.

The NAS was acting up – images were not being received on it. Had to reboot.



JD2455835

Observing log Posted on Oct 01, 2011 07:44AM

Rain on MLO. Moon up.

Tested setting the SKE from a script – this failed. The values for cusp offset and cusp angle were taken from the System_Dataxxx.csv master spreadsheet rather than the script itself (“protocol”).

We actually have no, documented, proof that the setting of the SKE or KEDFs works, right?



Dome and telescope pointing

Observing log Posted on Oct 01, 2011 07:21AM

Ben has created a map of where the dome should be for a range of telescope pointings. Map works well so far! We are adding data points to it as we go along, and also mapping out the edges of the dome slit at various positions:


A better resolution version is available here:

http://www.astro.utu.fi/~cflynn/dome_pointing_plot.ps



JD2455834

Observing log Posted on Sep 30, 2011 01:57PM

Testing the Focus Search tab in the Eng. Mode. Works nicely – but how do you reset the scale on the graph?

Also tried to ‘open’ and ‘close’ the Front Iris (elsewhere labelled as 30mm and Open). This produced no effect what so ever on diffraction spikes or flux levels. Does the Iris work? In which position is it stuck?



JD2455831

Observing log Posted on Sep 28, 2011 07:17PM

May have learned this important lesson:

The ‘flatlining’ of the CCD camera may be entirely due to the ‘running of two instances of LabView’. This occurs easily, it appears, if you close the VI for imaging while the Engineering Mode (EM) runs – or something like that: the danger is that an instance of Labview is started without you knowing it.
The danger is present when closing VIs. So, to guard against this, have the Task Manager running so that you can see whether you have more than one ‘instance of LabView’ running.

Also, you must have iBoot channel 5 ON when using any of the EM modes for controlling Filter Wheel, front Flap, Rear Flap, Side Flap, underside Flap or any other wonderful device on this telescope!



JD2455827

Observing log Posted on Sep 24, 2011 04:55PM

Sept 24 2011: Observed Jupiter with Chris. Did approx. 2000 0.1s V-band exposures. Used a script that contained direct GOTO RA/DEC command, called JUPITER.csv. Co-added the images on ‘woof’. Chris later analyzed the resulting profile.